• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 16:01:41 on September 17, 2005

    Entry number 154998

    keyword=Beam Time Accounting
    Hall A

    09/17/05 0800-1600
    Experiment #: E00-114
    Shift Leader: Mike Finn
    Shift Workers: Hachemi Benaoum

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    0.960.950.320.000.000.000.720.823.77
    Beam Available
    but Not in Use
    0.020.030.000.000.000.000.020.010.07
    Beam Not Available
    or Unacceptable
    0.020.030.681.001.001.000.260.174.17
    Accelerator Con-
    figuration Change
    0.000.000.000.000.000.000.000.000.00
    Experiment
    Ready
    1.001.001.001.001.001.001.001.008.00
    Planned Config-
    uration Change
    0.000.000.000.000.000.000.000.000.00
    Unplanned
    Experiment Down
    0.000.000.000.000.000.000.000.000.00
    Hall not
    ready reasons
    Target0.000.000.000.000.000.000.000.000.00
    Data Acq0.000.000.000.000.000.000.000.000.00
    Other0.000.000.000.000.000.000.000.000.00
    HRSL0.000.000.000.000.000.000.000.000.00
    HRSR0.000.000.000.000.000.000.000.000.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00

    Comments: 9:08 current had drifted up to 41uA request MCC lower it to 40uA 10:22 beam off for arc trip we are ending run and switching IHWP 10:42 IHWP set to IN positiion 11:17 asked MCC to reset Hall A sweeper magnet 13:52 MCC is ready to restore beam , ask for 2 uA to begin to check out spot size 14:17 beam back at 40uA resume production running

    Data Entry by: Mike Finn